# EW D-optimal Designs for Experiments with Mixed Factors

### v1

@@ Abstract
We consider EW D-optimal designs as robust designs for experiments under a general parametric model with discrete and continuous factors. When a pilot study is available, we recommend sample-based EW D-optimal designs for subsequent experiments. Otherwise, we recommend EW D-optimal designs under a prior distribution for model parameters. We propose an EW ForLion algorithm for finding EW D-optimal designs under a general parametric model, and justify that the designs found by our algorithm are EW D-optimal. To facilitate potential users in practice, we also develop a rounding algorithm that converts an approximate design with mixed factors to an exact design with prespecified grid points and the number of experimental units. By applying our algorithms for real experiments under multinomial logistic models or generalized linear models, we show that our designs are highly efficient with respect to locally D-optimal designs and more robust against parameter value misspecifications.
@@ 1 Introduction
In this paper, we consider robust designs for experiments with discrete and continuous factors. A motivating example is the paper feeder experiment described in Joseph and Wu (2004). The goal was to ensure precise feeding of one sheet of paper each time to a copier. Two common failure modes are frequently observed, namely misfeed when the feeder fails to feed any sheet of paper, and multifeed when the feeder picks up more than one sheets of paper at a time. The experiment involves eight discrete control factors (see Table 2 in Joseph and Wu (2004)) and one continuous variable, the stack force. The responses fall into three mutually exclusive categories, namely misfeed (typically with low stack force), normal, and multifeed (typically with high stack force). The results of a previous experiment were listed in Table 3 of Joseph and Wu (2004). The original design was conducted via a control array modified from $OA(18,2^{1}\times 3^{7})$ (see Table 5 in Joseph and Wu (2004)) for the eight discrete factors and an essentially uniform allocation on a predetermined set of discrete levels of the continuous factor. If a subsequent experiment will be conducted to study the factor effects, how can we do better?
@@
In the original analysis by Joseph and Wu (2004), two generalized linear models with probit link (McCullagh and Nelder, 1989; Dobson and Barnett, 2018) were used to model misfeed and multifeed separately. However, since misfeed and multifeed do not occur simultaneously and are both possible outcomes of the experiment, it would be more appropriate to adopt one multinomial model rather than two separate binomial models, so that the factor effects can be estimated more precisely due to the combined data and also be comparable for different types of failures. In this study, we adopt the multinomial logistic models (Glonek and McCullagh, 1995; Zocchi and Atkinson, 1999; Bu et al., 2020) for analyzing the paper feeder experiment, which include baseline-category (or known as multiclass logistic), cumulative, adjacent-categories, and continuation-ratio logit models.
@@
For experiments with discrete factors only, many numerical algorithms have been proposed for finding optimal designs, including Fedorov-Wynn (Fedorov, 1972; Fedorov and Hackl, 1997), multiplicative (Titterington, 1976, 1978; Silvey et al., 1978), cocktail (Yu, 2010), lift-one (Yang and Mandal, 2015; Yang et al., 2017; Bu et al., 2020), as well as classical optimization techniques (see Huang et al. (2024) for a review). For experiments with continuous factors only, Ai et al. (2023) incorporated the Fedorov-Wynn algorithm with the lift-one algorithm for continuation-ratio link models, Yang et al. (2013) and Harman et al. (2020) discretized the continuous factors into grid points and treated the experiments as with discrete factors.
@@
For experiments with mixed factors, particle swarm optimization (PSO) algorithms have been applied for D-optimal designs under generalized linear models with binary responses (Lukemire et al., 2019) and cumulative logit models (Lukemire et al., 2022), which, however, cannot guarantee the optimal solution to be ever found (Kennedy and Eberhart, 1995; Poli et al., 2007). Huang et al. (2024) proposed a ForLion algorithm, which added a merging step to Ai et al. (2023)’s algorithm and significantly reduced the number of design points. Unlike PSO-types of algorithms, the ForLion algorithm is guaranteed to find D-optimal designs based on the general equivalence theorem (Kiefer, 1974; Pukelsheim, 1993; Atkinson et al., 2007; Stufken and Yang, 2012; Fedorov and Leonov, 2014).
@@
The ForLion algorithm has been successfully applied to both generalized linear models and multinomial logistic models with mixed factors (Huang et al., 2024). However, it relies on pre-assumed parameter values and produces designs known as locally optimal (Chernoff, 1953), which may not be robust when the experimenters are uncertain about the parameter values. To address the limitation due to local optimality, Bayesian D-optimal designs have been proposed (Chaloner and Verdinelli, 1995), which maximize $E(\log|{\mathbf{F}}|)$ with respect to a prior distribution on unknown parameters, where ${\mathbf{F}}$ is the Fisher information matrix associated with the design. However, a major drawback of Bayesian D-optimal designs is its computational intensity, even with D-criterion that maximizes the determinant of the Fisher information matrix. As a promising alternative, Expected Weighted (EW) D-optimal designs have been proposed for generalized linear models (Yang et al., 2016), cumulative link models (Yang et al., 2017), and multinomial logistic models (Bu et al., 2020) with discrete factors, which maximize $\log|E({\mathbf{F}})|$ or equivalently $|E({\mathbf{F}})|$. Compared with Bayesian D-optimal designs, EW D-optimal designs are computationally much faster and often highly efficient in terms of the Bayesian criterion (Yang et al., 2016, 2017; Bu et al., 2020). Along this line, Tong et al. (2014) provided analytical solutions for certain cases of EW D-optimal designs for generalized linear models, and Bu et al. (2020) incorporated the lift-one and exchange algorithms with the EW criterion.
@@
In this paper, we adopt EW D-criterion and adapt the ForLion algorithm for finding robust designs with mixed factors. The original EW D-optimal designs were proposed for experiments with discrete factors or discretized continuous factors. To solve our problems, we formulate in Section 2 the EW D-optimal criterion and characterize EW D-optimal designs for general parametric statistical models with mixed factors. In Section 3, we develop algorithms for finding two types of EW D-optimal designs with mixed factors, namely sample-based EW D-optimal designs for experiments with a pilot study, and integral-based EW D-optimal designs for experiments with a prior distribution of unknown parameters. We further formulate EW D-optimal designs for multinomial logistic models and generalized linear models, and propose an algorithm for converting approximate designs with mixed factors to exact designs with grid points. In Section 4, we construct the proposed robust designs for two real experiments, namely the paper feeder experiment (Joseph and Wu, 2004), and an experiment on minimizing surface defects (Wu, 2008). In Section 5, we conclude our findings. The algorithms are implemented in R with version 4.4.2 and the simulations are run on a Windows 11 laptop with 32GB of RAM and a 13th Gen Intel Core i7-13700HX processor.
@@ 2 EW D-optimal Designs with Mixed Factors
Following Huang et al. (2024), we consider experiments with a design region ${\cal X}=\prod_{j=1}^{d}I_{j}\subset\mathbb{R}^{d}$, where $I_{j}$ is either a closed finite interval for the first $k$ factors, or a finite set of distinct numerical levels for $k+1\leq j\leq d$. In both cases, we denote $a_{j}=\min I_{j}>-\infty$ and $b_{j}=\max I_{j}<\infty$. Note that ${\cal X}$ here is compact, which is common for typical applications. If $k=0$, all factors are discrete or qualitative; if $k=d$, all factors are continuous.
@@
Given an experimental setting ${\mathbf{x}}_{i}=(x_{i1},\ldots,x_{id})^{T}\in{\cal X}$, suppose there are $n_{i}\geq 0$ experimental units assigned to this experimental setting. Their responses (could be vectors) are assumed to be iid from a parametric model $M({\mathbf{x}}_{i},\boldsymbol{\theta})$, where the parameter vector $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subseteq\mathbb{R}^{p}$ with the parameter space $\boldsymbol{\Theta}$. Suppose the responses are independent across different experimental settings. Under regularity conditions (see, e.g., Sections 2.5 and 2.6 in Lehmann and Casella (1998)), the corresponding Fisher information matrix can be written as $\sum_{i=1}^{m}n_{i}{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\theta})$, with respect to the corresponding design $\{({\mathbf{x}}_{i},n_{i}),i=1,\ldots,m\}$, known as an exact design, where ${\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\theta})$ is a $p\times p$ matrix, known as the Fisher information at ${\mathbf{x}}_{i}$ In design theory, $\boldsymbol{\xi}=\{({\mathbf{x}}_{i},w_{i}),i=1,\ldots,m\}$ with $w_{i}=n_{i}/n\geq 0$ and $n=\sum_{i=1}^{m}n_{i}$ , known as an approximate design, is often considered first (Kiefer, 1974; Pukelsheim, 1993; Atkinson et al., 2007), which is easier to find. The collection of all feasible designs is denoted by $\boldsymbol{\Xi}=\{\{({\mathbf{x}}_{i},w_{i}),i=1,\ldots,m\}\mid m\geq 1,{\mathbf{x}}_{i}\in{\cal X},w_{i}\geq 0,\sum_{i=1}^{m}w_{i}=1\}$.
@@ 2.1 EW D-optimal designs
Following Bu et al. (2020) and Huang et al. (2024), we adopt the D-criterion for optimal designs. When $\boldsymbol{\theta}$ is given or assumed to be known, the $\boldsymbol{\xi}\in\boldsymbol{\Xi}$ that maximizes $f(\boldsymbol{\xi})=|{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})|$ is known as a locally D-optimal approximate design (Chernoff, 1953), where ${\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})=\sum_{i=1}^{m}w_{i}{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\theta})$ is the Fisher information matrix associated with the design $\boldsymbol{\xi}$. The ForLion algorithm was proposed for finding locally D-optimal approximate designs with mixed factors (Huang et al., 2024).
@@
$f_{\rm EW}(\boldsymbol{\xi})=|E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}|=\left|\sum_{i=1}^{m}w_{i}E\left\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\right\}\right|\ ,$ (1)
@@
$E\left\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\right\}=\int_{\boldsymbol{\Theta}}{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\theta})Q(d\boldsymbol{\theta})$ (2)
is a $p\times p$ matrix after entry-wise expectation with respect to $Q(\cdot)$ on $\boldsymbol{\Theta}$. By replacing $\boldsymbol{\theta}$ in ${\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})$ with $\boldsymbol{\Theta}$, we indicate that the expectation $E\left\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\right\}$ is taken with respect to a random parameter vector.
When $\boldsymbol{\theta}$ is unknown but a dataset from a prior or pilot study is available, following Bu et al. (2020) and Huang et al. (2025), we bootstrap the original dataset to obtain a set of bootstrapped datasets, and the corresponding estimated parameter vectors $\{\hat{\boldsymbol{\theta}}_{1},\ldots,\hat{\boldsymbol{\theta}}_{B}\}$ by fitting the parametric model on each of the bootstrapped datasets. In other words, we may replace the prior distribution $Q(\cdot)$ in (2) with the empirical distribution of $\{\hat{\boldsymbol{\theta}}_{1},\ldots,\hat{\boldsymbol{\theta}}_{B}\}$. That is, we may look for $\boldsymbol{\xi}\in\boldsymbol{\Xi}$ maximizing
$f_{\rm SEW}(\boldsymbol{\xi})=|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}|=\left|\sum_{i=1}^{m}w_{i}\hat{E}\left\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\right\}\right|\ ,$ (3)
@@
$\hat{E}\left\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\right\}=\frac{1}{B}\sum_{j=1}^{B}{\mathbf{F}}({\mathbf{x}}_{i},\hat{\boldsymbol{\theta}}_{j})$ (4)
@@
For multinomial logistic models, such as the cumulative logit model used for the trauma clinical trial (see Example 5.2 in Bu et al. (2020)), the feasible parameter space $\boldsymbol{\Theta}=\{\boldsymbol{\theta}=(\beta_{11},\beta_{12},\beta_{21},\beta_{22},\beta_{31},\beta_{32},\beta_{41},\beta_{42})^{T}\mid\beta_{11}+\beta_{12}x<\beta_{21}+\beta_{22}x<\beta_{31}+\beta_{32}x<\beta_{41}+\beta_{42}x,\mbox{ for all }x\in{\mathcal{X}}\}$ is not rectangular, which makes the integral in (2) difficult to handle. In this case, even with a prior distribution instead of a dataset, we may also draw a random sample $\{\hat{\boldsymbol{\theta}}_{1},\ldots,\hat{\boldsymbol{\theta}}_{B}\}$ from the prior distribution, and adopt the sample-based EW D-optimality. According to Section S6 in the Supplementary Material, in terms of relative efficiency, the obtained EW D-optimal designs are fairly stable against the random samples.

### v2

@@ Abstract
We characterize EW D-optimal designs as robust designs against unknown parameter values for experiments under a general parametric model with discrete and continuous factors. When a pilot study is available, we recommend sample-based EW D-optimal designs for subsequent experiments. Otherwise, we recommend EW D-optimal designs under a prior distribution for model parameters. We propose an EW ForLion algorithm for finding EW D-optimal designs with mixed factors, and justify that the designs found by our algorithm are EW D-optimal. To facilitate potential users in practice, we also develop a rounding algorithm that converts an approximate design with mixed factors to exact designs with prespecified grid points and the number of experimental units. By applying our algorithms for real experiments under multinomial logistic models or generalized linear models, we show that our designs are highly efficient with respect to locally D-optimal designs and more robust against parameter value misspecifications.
@@ 1 Introduction
In this paper, we consider robust designs against unknown parameter values for experiments with discrete and continuous factors. A motivating example is the paper feeder experiment described in Joseph and Wu (2004). The goal was to ensure precise feeding of one sheet of paper each time to a copier. Two common failure modes are frequently observed, namely misfeed when the feeder fails to feed any sheet of paper, and multifeed when the feeder picks up more than one sheets of paper at a time. The experiment involves eight discrete control factors (see Table 2 in Joseph and Wu (2004) and also Section S4 in the Supplementary Material) and one continuous factor, the stack force. The responses fall into three mutually exclusive categories, namely misfeed (typically with low stack force), normal, and multifeed (typically with high stack force). The results of the original experiment were listed in Table 3 of Joseph and Wu (2004). The original design was conducted via a control array modified from $OA(18,2^{1}\times 3^{7})$ (see Table 5 in Joseph and Wu (2004)) for the eight discrete factors and an essentially uniform allocation on a predetermined set of discrete levels of the continuous factor. If a subsequent experiment will be conducted to study the factor effects, how can we do better?
@@
In the original analysis by Joseph and Wu (2004), two generalized linear models (McCullagh and Nelder, 1989; Dobson and Barnett, 2018) with probit link were used to model misfeed and multifeed separately. However, since misfeed and multifeed do not occur simultaneously and are both possible outcomes of the experiment, it would be more appropriate to adopt one multinomial model rather than two separate binomial models, so that the factor effects can be estimated more precisely due to the combined data and also be comparable for different types of failures. In this study, we adopt the multinomial logistic models (Glonek and McCullagh, 1995; Zocchi and Atkinson, 1999; Bu et al., 2020) for analyzing the paper feeder experiment, which include baseline-category (also known as multiclass logistic), cumulative, adjacent-categories, and continuation-ratio logit models.
@@
For experiments with discrete factors only, many numerical algorithms have been proposed for finding optimal designs, including Fedorov-Wynn (Fedorov, 1972; Fedorov and Hackl, 1997), multiplicative (Titterington, 1976, 1978; Silvey et al., 1978), cocktail (Yu, 2010), lift-one (Yang and Mandal, 2015; Yang et al., 2017; Bu et al., 2020), as well as classical optimization techniques (see Huang et al. (2024) for a review). According to Yang et al. (2016) and Huang et al. (2024), the lift-one algorithm outperforms commonly used optimization algorithms and obtains D-optimal designs with fewer distinct design points. For experiments with continuous factors only, Ai et al. (2023) incorporated the Fedorov-Wynn algorithm with the lift-one algorithm for continuation-ratio link models, Yang et al. (2013) and Harman et al. (2020) discretized the continuous factors into grid points and treated the experiments as with discrete factors.
@@
For experiments with mixed factors, particle swarm optimization (PSO) algorithms have been applied for D-optimal designs under generalized linear models with binary responses (Lukemire et al., 2019) and cumulative logit models (Lukemire et al., 2022), which, however, cannot guarantee the optimal solution to be ever found (Kennedy and Eberhart, 1995; Poli et al., 2007). Huang et al. (2024) proposed a ForLion algorithm, which added a merging step to Ai et al. (2023)’s algorithm and significantly reduced the number of design points. In practice, it often leads to reduced experimental time and cost (Huang et al., 2024). Unlike PSO-types of algorithms, the ForLion algorithm is guaranteed to find D-optimal designs based on the general equivalence theorem (Kiefer, 1974; Pukelsheim, 1993; Atkinson et al., 2007; Stufken and Yang, 2012; Fedorov and Leonov, 2014).
@@
The ForLion algorithm has been successfully applied to both generalized linear models and multinomial logistic models with mixed factors (Huang et al., 2024). However, it relies on pre-assumed parameter values and produces designs known as locally optimal (Chernoff, 1953), which may not be robust when the parameter values are misspecified. To address the limitation due to local optimality, Bayesian D-optimal designs have been proposed (Chaloner and Verdinelli, 1995), which maximize $E(\log|{\mathbf{F}}|)$ with respect to a prior distribution on unknown parameters, where ${\mathbf{F}}$ is the Fisher information matrix associated with the design. However, a major drawback of Bayesian D-optimal designs is its computational intensity, even with D-criterion that maximizes the determinant of the Fisher information matrix. As a promising alternative, Expected Weighted (EW) D-optimal designs have been proposed for generalized linear models (Yang et al., 2016), cumulative link models (Yang et al., 2017), and multinomial logistic models (Bu et al., 2020) with discrete factors, which maximize $\log|E({\mathbf{F}})|$ or equivalently $|E({\mathbf{F}})|$. Compared with Bayesian D-optimal designs, EW D-optimal designs are computationally much faster and often highly efficient in terms of the Bayesian criterion (Yang et al., 2016, 2017; Bu et al., 2020). Along this line, Tong et al. (2014) provided analytical solutions for certain cases of EW D-optimal designs for generalized linear models, and Bu et al. (2020) incorporated the lift-one and exchange algorithms with the EW criterion.
@@
In this paper, we adopt EW D-criterion and adapt the ForLion algorithm for finding robust designs with mixed factors. The original EW D-optimal designs were proposed for experiments with discrete factors or discretized continuous factors. To solve our problems, we formulate in Section 2 the EW D-optimal criterion and characterize EW D-optimal designs for general parametric statistical models with mixed factors. In Section 3, we develop algorithms for finding two types of EW D-optimal designs with mixed factors, namely sample-based EW D-optimal designs for experiments with a pilot study, and integral-based EW D-optimal designs for experiments with a prior distribution for unknown parameters. We further formulate EW D-optimal designs for multinomial logistic models and generalized linear models, and propose an algorithm for converting approximate designs with mixed factors to exact designs with grid points, which is much faster than finding optimal designs restricted on the same set of grid points. In Section 4, we construct the proposed robust designs for two real experiments, namely the paper feeder experiment (Joseph and Wu, 2004), and an experiment on minimizing surface defects (Wu, 2008). In Section 5, we conclude our findings. The algorithms are implemented in R with version 4.4.2 and the simulations are run on a Windows 11 laptop with 32GB of RAM and a 13th Gen Intel Core i7-13700HX processor.
@@ 2 EW D-optimal Designs with Mixed Factors
Following Huang et al. (2024), we first consider experiments with a design region ${\cal X}=\prod_{j=1}^{d}I_{j}\subset\mathbb{R}^{d}$, where $I_{j}$ is either a closed finite interval for the first $k$ factors, or a finite set of distinct numerical levels for $k+1\leq j\leq d$. In both cases, we denote $a_{j}=\min I_{j}>-\infty$ and $b_{j}=\max I_{j}<\infty$. If $k=0$, all factors are discrete or qualitative; if $k=d$, all factors are continuous. Note that ${\cal X}$ here is compact and built as a product of sets, which is common for typical applications.
For some applications, for examples, when the levels of some discrete factors are not numerical (see Example 1 below), or the level combinations of discrete factors under consideration are restricted to a true subset of all possible level combinations (see Example 2 below), we also need to consider design regions ${\cal X}=\prod_{j=1}^{k}I_{j}\ \times\ {\cal D}$, where ${\cal D}\subseteq\prod_{j=k+1}^{d}I_{j}$ Apparently, $\prod_{j=1}^{d}I_{j}$ is a special case of $\prod_{j=1}^{k}I_{j}\ \times\ {\cal D}$ with ${\cal D}=\prod_{j=k+1}^{d}I_{j}$
Example 1.
A polysilicon deposition process for circuit manufacturing was described by Phadke (1989) involved one qualitative control factor, cleaning method (None for no cleaning, $CM_{2}$ for cleaning inside the reactor, and $CM_{3}$ for cleaning outside the reactor), and five other factors that are continuous in nature (see Table 4.1 in Phadke (1989)). Since the levels of cleaning method are not numerical, it would be more appropriate to convert it to two (binary) indicator variables $({\mathbf{1}}_{CM_{2}},{\mathbf{1}}_{CM_{3}})$ with $(0,0),(1,0),(0,1)$ standing for None, $CM_{2}$, and $CM_{3}$, respectively. In this case, the corresponding design region is ${\cal X}=\prod_{j=1}^{5}I_{j}\ \times\ \{(0,0),(1,0),(0,1)\}$, where $I_{1},\ldots,I_{5}$ may be defined as in Table S1 of the Supplementary Material of Huang et al. (2024). Note that the cleaning method was simplified into a binary factor ($CM_{2}$ or $CM_{3}$) in Lukemire et al. (2022) and Huang et al. (2024). $\Box$
Example 2.
In the paper feeder experiment described by Joseph and Wu (2004), besides one continuous factor, there are eight discrete or qualitative control factors, including two binary ones and six three-level ones (see Section S4 in the Supplementary Material). Instead of considering all $9162^{2}\times 3^{6}=2,$ 916 level combinations, the experimenters adopted 18 out of them modified from an orthogonal array $OA(18,2^{1}\times 3^{7})$ (see Table 5 in Joseph and Wu (2004)). In this case, for illustration purposes, we may consider a restricted design region ${\cal X}=[0,160]\times{\cal D}$, where $[0,160]$ is the range of the continuous factor, and ${\cal D}=\{(x_{i2},\ldots,x_{i9})\mid i=1,\ldots,18\}$ consists of the original 18 level combinations of eight discrete factors. $\Box$
@@
Given an experimental setting ${\mathbf{x}}_{i}=(x_{i1},\ldots,x_{id})^{T}\in{\cal X}$, suppose there are $n_{i}\geq 0$ experimental units assigned to this experimental setting. Their responses (could be vectors) are assumed to be iid from a parametric model $M({\mathbf{x}}_{i},\boldsymbol{\theta})$, where the parameter vector $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subseteq\mathbb{R}^{p}$ with the parameter space $\boldsymbol{\Theta}$. Suppose the responses are independent across different experimental settings. Under regularity conditions (see, e.g., Sections 2.5 and 2.6 in Lehmann and Casella (1998)), the corresponding Fisher information matrix can be written as $\sum_{i=1}^{m}n_{i}{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\theta})$, with respect to the corresponding design $\{({\mathbf{x}}_{i},n_{i}),i=1,\ldots,m\}$, known as an exact design, where ${\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\theta})$ is a $p\times p$ matrix, known as the Fisher information at ${\mathbf{x}}_{i}$ In design theory, $\boldsymbol{\xi}=\{({\mathbf{x}}_{i},w_{i}),i=1,\ldots,m\}$ with $w_{i}=n_{i}/n\geq 0$ and $n=\sum_{i=1}^{m}n_{i}$ , known as an approximate design, is often considered first (Kiefer, 1974; Pukelsheim, 1993; Atkinson et al., 2007). The collection of all feasible designs is denoted by $\boldsymbol{\Xi}=\{\{({\mathbf{x}}_{i},w_{i}),i=1,\ldots,m\}\mid m\geq 1,{\mathbf{x}}_{i}\in{\cal X},w_{i}\geq 0,\sum_{i=1}^{m}w_{i}=1\}$.
@@ 2.1 EW D-optimal designs
Following Bu et al. (2020) and Huang et al. (2024), we adopt the D-criterion for optimal designs. When $\boldsymbol{\theta}$ is given or assumed to be known, an $\boldsymbol{\xi}\in\boldsymbol{\Xi}$ that maximizes $f(\boldsymbol{\xi})=|{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})|$ is known as a locally D-optimal design (Chernoff, 1953), where ${\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})=\sum_{i=1}^{m}w_{i}{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\theta})$ is the Fisher information matrix associated with $\boldsymbol{\xi}$. The ForLion algorithm was proposed by Huang et al. (2024) for finding locally D-optimal designs with mixed factors.
@@
$f_{\rm EW}(\boldsymbol{\xi})=|E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}|=\left|\sum_{i=1}^{m}w_{i}E\left\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\right\}\right|\ ,$ (2.1)
@@
$E\left\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\right\}=\int_{\boldsymbol{\Theta}}{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\theta})Q(d\boldsymbol{\theta})$ (2.2)
is a $p\times p$ matrix after entry-wise expectation with respect to $Q(\cdot)$ on $\boldsymbol{\Theta}$. By replacing $\boldsymbol{\theta}$ in ${\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})$ with $\boldsymbol{\Theta}$, we indicate that the expectation $E\left\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\right\}$ is taken with respect to a random parameter vector in $\boldsymbol{\Theta}$.
When $\boldsymbol{\theta}$ is unknown but a dataset from a prior or pilot study is available, following Bu et al. (2020) and Huang et al. (2025), we bootstrap the original dataset to obtain a set of bootstrapped datasets, and their corresponding parameter vectors $\{\hat{\boldsymbol{\theta}}_{1},\ldots,\hat{\boldsymbol{\theta}}_{B}\}$ obtained by fitting the parametric model on each of the bootstrapped datasets. In other words, we may replace the prior distribution $Q(\cdot)$ in (2.2) with the empirical distribution of $\{\hat{\boldsymbol{\theta}}_{1},\ldots,\hat{\boldsymbol{\theta}}_{B}\}$. That is, we may look for $\boldsymbol{\xi}\in\boldsymbol{\Xi}$ maximizing
$f_{\rm SEW}(\boldsymbol{\xi})=|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}|=\left|\sum_{i=1}^{m}w_{i}\hat{E}\left\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\right\}\right|\ ,$ (2.3)
@@
$\hat{E}\left\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\right\}=\frac{1}{B}\sum_{j=1}^{B}{\mathbf{F}}({\mathbf{x}}_{i},\hat{\boldsymbol{\theta}}_{j})$ (2.4)
@@
For multinomial logistic models, such as the cumulative logit model used for the trauma clinical trial (see Example 5.2 in Bu et al. (2020)), the feasible parameter space $\boldsymbol{\Theta}=\{\boldsymbol{\theta}=(\beta_{11},\beta_{12},\beta_{21},\beta_{22},\beta_{31},\beta_{32},\beta_{41},\beta_{42})^{T}\mid\beta_{11}+\beta_{12}x<\beta_{21}+\beta_{22}x<\beta_{31}+\beta_{32}x<\beta_{41}+\beta_{42}x,\mbox{ for all }x\in{\mathcal{X}}\}$ is not rectangular, which makes the integral in (2.2) difficult to calculate. In this case, even with a prior distribution instead of a dataset, we may also draw a random sample $\{\hat{\boldsymbol{\theta}}_{1},\ldots,\hat{\boldsymbol{\theta}}_{B}\}$ from the prior distribution, and adopt the sample-based EW D-optimality. According to Section S5 in the Supplementary Material, in terms of relative efficiency, the obtained EW D-optimal designs are fairly stable against the random samples.